# Impact of Calibration and Position Errors on Astrophysical Parameters of the Hi 21cm Signal

### v2

@@ Abstract
The Epoch of Reionization (EoR) and Cosmic Dawn (CD) are pivotal stages during the first billion years of the universe, exerting a significant influence on the development of cosmic structure. The detection of the redshifted 21-cm signal from these epochs is challenging due to the dominance of significantly stronger astrophysical foregrounds and the presence systematics. This work used the 21cm E2E (end to end) pipeline, followed by simulation methodology described [1] to conduct synthetic observations of a simulated sky model that includes both the redshifted 21-cm signal and foregrounds. A framework was constructed using Artificial Neural Networks (ANN) and Bayesian techniques to directly deduce astrophysical parameters from the measured power spectrum. This approach eliminates the need for explicit telescope layout effects correction in interferometric arrays such as SKA-Low. The present work investigates the impact of gain calibration and sky model position errors on the recovery of the redshifted 21-cm power spectrum for the SKA-Low AA array configuration. We assessed the effects of these inaccuracies on the deduced astrophysical parameters and established acceptable tolerance levels. Based on our results, the gain calibration error tolerance for ideal signal detection is 0.001 %. However, if the sky model position errors exceed 0.048 arcseconds, the remaining foregrounds would obscure the target signal.
@@ 1 Introduction
The redshifted 21cm line is a potential probe of the early Universe [2, 3, 4], mainly from the era post recombination until the universe became fully ionized. Based on the theoretical model, in the first billion years of the Universe, Cosmic Dawn (CD) is when the first star or galaxy will be formed (30 $>z>$ 12). These stars and galaxies are formed due to gravitational instability, which causes small-scale fluctuation in the matter density. The UV photons produced by these objects ionized the neutral hydrogen (Hi) in the intergalactic medium (IGM). This transition period is known as Epoch of Reionization (EoR) [5, 6, 7, 8, 9]. Based on various indirect observations such as quasar absorption [10] at high redshifts, Thompson scattering optical depth [11] suggest that the reionization extended the process and lasted at least til the redshift $z\sim 6$.
The observation of Hi 21cm power spectrum using large interferometric arrays currently holds the most significant potential to observe the redshifted Hi 21cm line [12, 13, 14, 15]. The 21cm signal inherently encodes information about the underlying dark matter distribution and the properties of the ionizing sources. As a result, it has the potential to trace the history of reionization, reflecting the evolution of the average ionization state of the intergalactic medium (IGM) with redshift during the EoR. Detection of the redshifted Hi signal is the key science goal of several ongoing and future experiments. Recently, the detection of a global 21cm signal reported by the Experiment to Detect Global Epoch of Reionization Signature (EDGES) team [16]. However, the detection has been challenged by another independent experiment, SARAS [17, 18]. Besides EDGES and SARAS, there are other independent single antenna experiments like BIGHORNS [19], SCI-HI [20], and LEDA [21] are also aiming to detect the global signal but are yet to report any detection. Conversely, interferometers focusing on statistical fluctuations have yielded significant upper limits on the EoR power spectrum (PS) amplitude. The most sensitive operational interferometers, such as the GMRT, MWA, LOFAR, and HERA, have all established upper limits on the power spectrum amplitude of the signal [22, 23, 24, 25, 26], but there has been no confirmed detection of the cosmological Hi 21cm signal.
@@
[…]
@@ 2.1 Hi 21cm Maps
We simulated lightcones in a 500500500 $Mpc^{3}$ comoving box with $232^{3}$ of grid cells for a range for redshift extent of 8.738.738.73 $\leq$ $z$leq 9.299.299.29. The constructed lightcone volume is saved on a grid of size 232 × 232 × 64 and projected on a World Coordinate System (WCS) to generate an input signal for simulating observations. For this study, we simulate different lightcones to probe different realizations of reionization history using varying EoR parameters. To simulate these lightcones, we use three key EoR parameters $R_{\mathrm{mfp}}$, $T_{\mathrm{vir}}$, and $\zeta$ which can be tweaked to create different reionization histories. These parameters are commonly used to characterize the EoR, as they effectively capture the timing, source population, and morphology of reionization while remaining physically interpretable and computationally efficient [85]. We describe these in detail below following [85, 76]:
@@
  • $\rm T_{\mathrm{vir}}$, Minimum virial temperature of haloes producing ionizing photons: This represents the minimum mass of haloes producing ionizing photons during the EoR. Usually, $\rm T_{\mathrm{vir}}$ is chosen to be larger than $10^{4}$ such that atomic cooling become effective [88, 89].
  • $\rm\zeta$, Ionizing efficiency: Ionizing efficiency refers to the ability of sources, such as stars or galaxies, to convert their energy into ionizing photons that can ionize hydrogen in the intergalactic medium. This is a combination of several degenerate astrophysical parameters and is defined as $\zeta$ = $f_{\mathrm{esc}}$$f_{\mathrm{\star}}$ $N_{\mathrm{\gamma}}$/(1 + $n_{\mathrm{rec}})$ [90, 91]. Here, $\rm f_{esc}$ is the fraction of ionizing photons escaping from galaxies into the IGM, $\rm f_{\star}$ is the fraction of baryons locked into stars and $\rm N_{\mathrm{\gamma}}$ is the number of ionizing photons produced per baryon in stars and $\rm n_{\mathrm{rec}}$ is the mean recombination rate per baryon.
@@ 3 Synthetic Observations
To conduct synthetic observations, we utilized a 21cm end-to-end (E2E) pipeline [1]. This pipeline employs the OSKAR software package [93] for simulating SKA-Low configurations, the Common Astronomy Software Applications (CASA) package [94] for further reading and processing of the visibility data. In this simulated observation, the sky was observed with a phase center at $\alpha=15^{\text{h}}00^{\text{m}}00^{\text{s}}$ and $\delta=-30^{\circ}$ for a duration of 4 hours (±2$\pm$ 2± 2 hours hour angle). The observing bandwidth of the lightcone spans 8 MHz with a channel separation of 125 kHz. For additional details, refer to Tab. 1.
@@ 3.1 Telescope Model
Figure 1: Telescope configurations utilized in the simulation, shown from left to right: SKA-Low AA* (with a 2 km central core).
@@ 4.1 Theoretical Power Spectrum
We can define the dimension 21cm power spectrum as :
@@ 4.2 Observational Power Spectrum
[…]
@@ 5 Emulating Hi 21cm Power Spectrum
Figure 3: The figure shows a comparison between the simulated true power spectrum (solid lines) and the emulated power spectrum by the ANN (dots). The left panel shows emulator predictions trained on theoretical power spectra, while the right panel presents predictions based on observed power spectra with SKA-Low.
@@ 7 Error Covariances of Power Spectrum
where $\rm A_{\mathrm{eff}}$, $\rm A_{\mathrm{coll}}$, and $\rm A_{\mathrm{core}}$ denote the effective area, collecting area, and core area, respectively, of the specified radio interferometric array. The symbol $\rm t_{\mathrm{obs}}$ signifies the total observation time in hours. Additionally, $\rm D$ and $\rm\Delta D$ denote the comoving distance to the redshifts where the center is located and the comoving distance corresponding to a bandwidth $\rm B$ at that comoving distance. $\rm T_{sky}$ denotes the sky temperature calculated using $\mathrm{T_{\mathrm{sky}}\sim 180\Big{(}\frac{\nu}{180MHz}\Big{)}^{-2.6}}$ [5], symbol $\Omega_{\mathrm{FoV}}$ denotes field of view symbol, $n_{\mathrm{p}}$ represents the number of polarization and $N_{\mathrm{k_{i}}}$ denotes the number of measurement at each k-mode.
@@ 8.1 Perfect Observing Condition
We also noticed that the power spectrum from set 1, whose astrophysical parameters are around the centre of the boundaries, generated more accurate conclusions. By contrast, the power spectra generated in close proximity to the boundary values of the astrophysical parameters yielded inferred parameter mean that slightly differed from the actual value. For example, in Set 3, the true value of $\rm\zeta$ is 57.42, while the inferred mean using the theoretical power spectrum is 50.8. Using the SKA power spectrum, the inferred mean is 59.8. Although these values differ from the true mean, they still lie within $1\sigma$ of the inferred uncertainty range (see Table 2). This mismatch could be due to the ANN emulator being trained on less training datasets. Potential future development of a more precise emulation could involve training it with a more extensive dataset that cover the parameter space in the effective manner.
@@ 8.2.1 Gain Calibration Errors
To assess the influence of gain calibration errors on the observed power spectrum (PS) and their subsequent impact on associated astrophysical parameters for SKA-Low, we introduced various gain calibration errors as detailed in [1]. The resulting residual PS was visually compared with the actual PS, depicted in the top left panel of Fig. 5. The parameter obtained from the PS analysis with SKA-Low, incorporating actual signal and PS for gain error residuals, is depicted in the top right panel of Fig. 5. Upon examining the plotted residual PS spectrum, we observed that for gain calibration error of 0.001 %, the residual PS overlaps with the true PS. For residual PS with a gain calibration error of 0.01%, we observe overlapping at lower k modes but with minor deviation at higher k modes. However, when deriving astrophysical parameters from the residual power spectrum, we found that for a gain calibration error of 0.001%, the derived parameters match those derived from the true PS. In contrast, for PS with the gain calibration error of 0.01%, the derived parameters exhibit significant differences from the actual values. This indicates that astrophysical parameters are highly sensitive to gain calibration errors, and even minor deviations in the PS due to gain calibration errors can severely impact the estimation of parameters, introducing bias. The sensitivity of derived parameters to gain calibration errors clearly suggests that the gain calibration error threshold should be much lower than 0.01%, ideally closer to 0.001%. If the gain calibration error exceeds 0.001%, it is advisable to model them to prevent bias in the derived parameter values.
@@ 8.2.2 Sky Model Position Error
Similar to gain calibration errors, we evaluated the impact of sky model position errors on the observed power spectrum (PS) and their effect on the inferred astrophysical parameters for SKA-Low. Different sky model position errors, as described above, were introduced, and the resulting residual PS was compared with the true PS, as shown in the top-left panel of Fig. 6. The parameters obtained from the PS, which incorporates actual signals and position error residuals, are depicted in the top right panel of Fig. 6. Upon examining the plotted residual PS spectrum, we found that for a sky model position error of 0.048”, the residual PS overlaps with the true PS. At lower k modes, the residual PS with this position error overlaps well with the true PS, but at higher k modes, there is a minor deviation. This occurs because real-space scales are more affected by arcsecond-level errors at higher k values. However, since we used the entire k range to derive astrophysical parameters from the residual power spectrum, we observed that for a position error of 0.048”, the derived parameters closely match those obtained from the true PS. Similarly, for a sky model position error of 0.48”, the derived parameters exhibit substantial deviations from the true values. This indicates that astrophysical parameters are highly sensitive to sky model position errors, and even small deviations in the PS due to these errors can severely impact parameter estimation, introducing bias. The sensitivity of derived parameters to sky model position errors, similar to gain calibration errors, suggests that the acceptable sky model position error threshold should be much lower than 0.48”, ideally around 0.048”. If the sky model position error exceeds 0.048”, modeling the errors becomes crucial to prevent bias in the derived parameter values.
@@ 9 Summary and Discussion
This study presents the development of an ANN and Bayesian-based framework for inferring astrophysical parameters from observed Hi power spectra (PS). We used the 21cm E2E pipeline to generate the mock observed power spectrum (PS) for SKA-Low core array configurations. However, generating a large number of modelled observable signals across a multi-dimensional parameter space for Bayesian inference is computationally expensive, particularly when using simulated observations. To address this, we adopted a formalism utilized by several groups, employing emulators for modelling the EoR signal instead of actual simulations.
@@
  • The position error threshold for SKA-Low is found to be <0.48<0.48"< 0.48 ”, as displacements >0.048>0.048"> 0.048 ” cause significant deviations between the inferred astrophysical parameters and their actual values. This suggests that sky model position errors can bias the inference of the astrophysical parameters.
@@
Thus, we conclude that for the existing interferometers, the gain calibration error should be nearly 0.001%, and the sky model position error should be less than 0.48” to avoid biasing the inference. Beyond these thresholds, these errors start to affect the results. To reduce their impact, calibration must be performed with the same level of accuracy, or we need to develop modeling or mitigation techniques to eliminate the residual foregrounds arising from calibration or sky model inaccuracies, ensuring minimal influence on the inference. It is important to note that the simulations presented in this study rely on several simplifying assumptions regarding foregrounds and instrumental effects. This underscores the need for more realistic and detailed simulations, including factors such as primary beam chromaticity and other instrumental systematics, to more accurately assess the robustness of parameter inference and to inform optimal array design. These complexities will be addressed in future work for a more comprehensive understanding.

### v3

@@ Abstract
The Epoch of Reionization (EoR) and Cosmic Dawn (CD) are pivotal stages during the first billion years of the universe, exerting a significant influence on the development of cosmic structure. The detection of the redshifted 21-cm signal from these epochs is challenging due to the dominance of significantly stronger astrophysical foregrounds and the presence of systematics. This work used the 21cm E2E (end to end) pipeline, followed by simulation methodology described [1] to conduct synthetic observations of a simulated sky model that includes both the redshifted 21-cm signal and foregrounds. A framework was constructed using Artificial Neural Networks (ANN) and Bayesian techniques to directly deduce astrophysical parameters from the measured power spectrum. This approach eliminates the need for explicit telescope layout effects correction in interferometric arrays such as SKA-Low. The present work investigates the impact of gain calibration errors and sky model position errors on the recovery of the redshifted 21-cm power spectrum for the SKA-Low $A$ array configuration. We assessed the effects of these inaccuracies on the deduced astrophysical parameters and established acceptable tolerance levels. Based on our results, the gain calibration error tolerance for ideal signal detection is 0.001 %. However, if the sky model position errors exceed 0.048 arcseconds, the remaining foregrounds would obscure the target signal.
@@ 1 Introduction
The redshifted 21cm line is a potential probe of the early Universe [2, 3, 4], mainly from the era post recombination until the universe became fully ionized. Based on the theoretical model, in the first billion years of the Universe, Cosmic Dawn (CD) is when the first star or galaxy will be formed (30 >z>>z> 12). These stars and galaxies are formed due to gravitational instability, which causes small-scale fluctuation in the matter density. The UV photons produced by these objects ionized the neutral hydrogen (Hi) in the intergalactic medium (IGM). This transition period is known as Epoch of Reionization (EoR) [5, 6, 7, 8, 9]. Based on various indirect observations such as quasar absorption [10] at high redshifts, Thompson scattering optical depth [11] suggest that the reionization extended the process and lasted at least til the redshift $z\sim 6$.
The observation of Hi 21cm power spectrum using large interferometric arrays currently holds the most significant potential to observe the redshifted Hi 21cm line [12, 13, 14, 15]. The 21cm signal inherently encodes information about the underlying dark matter distribution and the properties of the ionizing sources. As a result, it has the potential to trace the history of reionization, reflecting the evolution of the average ionization state of the intergalactic medium (IGM) with redshift during the EoR. Detection of the redshifted Hi signal is the key science goal of several ongoing and future experiments. Recently, the detection of a global 21cm signal reported by the Experiment to Detect Global Epoch of Reionization Signature (EDGES) team [16]. However, the detection has been challenged by another independent experiment, SARAS [17, 18]. Besides EDGES and SARAS, there are other independent single radiometer experiments like BIGHORNS [19], SCI-HI [20], and LEDA [21] are also aiming to detect the global signal but have not yet to report any detection. Conversely, interferometers focusing on statistical fluctuations have yielded significant upper limits on the EoR power spectrum (PS) amplitude. The most sensitive operational interferometers, such as the GMRT, MWA, LOFAR, and HERA, have all established upper limits on the power spectrum amplitude of the signal [22, 23, 24, 25, 26], but there has been no confirmed detection of the cosmological Hi 21cm signal.
@@
The signal-to-noise ratio (SNR) is a critical factor in the detection of faint cosmic signals, such as the redshifted 21-cm line from the epoch of reionization. [34] demonstrated that even minor calibration errors can significantly reduce the dynamic range, thereby obscuring the cosmological signal. Building on this, [1], through an analysis of the one-dimensional power spectrum (1D PS), determined that the optimal calibration error tolerance for reliable signal detection is approximately $0.01\%$. Consistently, [63] showed that calibration errors must be limited to below approximately $10^{-5}$ or 0.001 % in amplitude in order to prevent contamination of the EoR window in power spectrum measurements. Despite these insights, further investigation is needed to determine the precise tolerance limits for various instrumental imperfections that could hinder weak signal detection. This study aims to quantify the tolerance levels required for the successful detection of the redshifted 21-cm signal and the recovery of astrophysical parameters from the epoch of reionization using the SKA-Low telescope. To achieve this, we utilize a hybrid machine learning (ML) approach that integrates artificial neural networks (ANN) with Bayesian methods to analyze the effects of imperfections on observational data and their direct impact on the associated astrophysical parameters.
@@ 2.1 Hi 21cm Maps
We simulated lightcones in a 500500 $Mpc^{3}$ comoving box with $232^{3}$ of grid cells for a range for redshift extent of $8.73$ $\leq$ $z$ $\leq$ $9.29$. The constructed lightcone volume is saved on a grid of size 232 × 232 × 64 and projected on a World Coordinate System (WCS) to generate an input signal for simulating observations. For this study, we simulate different lightcones to probe different realizations of reionization history using varying EoR parameters. To simulate these lightcones, we use three key EoR parameters $R_{\mathrm{mfp}}$, $T_{\mathrm{vir}}$, and $\zeta$ which can be tweaked to create different reionization histories. These parameters are commonly used to characterize the EoR, as they effectively capture the timing, source population, and morphology of reionization while remaining physically interpretable and computationally efficient [85]. We describe these in detail below following [85, 76]:
@@
  • $\rm T_{\mathrm{vir}}$, Minimum virial temperature of haloes producing ionizing photons: This represents the minimum mass of haloes producing ionizing photons during the EoR. Usually, $\rm T_{\mathrm{vir}}$ is chosen to be larger than $10^{4}$K such that atomic cooling become effective [88, 89].
  • $\rm\zeta$, Ionizing efficiency: Ionizing efficiency refers to the ability of sources, such as stars or galaxies, to convert their energy into ionizing photons that can ionize hydrogen in the intergalactic medium. This is a combination of several degenerate astrophysical parameters and is defined as $\zeta$ = $_{\mathrm{esc}}ff_{\mathrm{\star}}$ $N_{\mathrm{\gamma}}$/(1 + $n_{\mathrm{rec}})$ [90, 91]. Here, $\rm f_{esc}$ is the fraction of ionizing photons escaping from galaxies into the IGM, $\rm f_{\star}$ is the fraction of baryons locked into stars and $\rm N_{\mathrm{\gamma}}$ is the number of ionizing photons produced per baryon in stars and $\rm n_{\mathrm{rec}}$ is the mean recombination rate per baryon.
@@ 3 Synthetic Observations
To conduct synthetic observations, we utilized a 21cm end-to-end (E2E) pipeline [1]. This pipeline employs the OSKAR software package [93] for simulating SKA-Low configurations, the Common Astronomy Software Applications (CASA) package [94] for further reading and processing of the visibility data. In this simulated observation, the sky was observed with a phase center at $\alpha=15^{\text{h}}00^{\text{m}}00^{\text{s}}$ and $\delta=-30^{\circ}$ for a duration of 4 hours ($\pm 2$ hours hour angle). The observing bandwidth of the lightcone spans 8 MHz with a channel separation of 125 kHz. For additional details, refer to Tab. 1.
@@ 3.1 Telescope Model
Figure 1: Telescope configuration utilized in the simulation: SKA-Low AA* (with a 2 km central core).
@@ 4.1 Theoretical Power Spectrum
We can define the dimensionless 21cm power spectrum as :
@@ 4.2 Observational Power Spectrum
$V({\mathbf{U}},\nu)=\iint A(\hat{\mathbf{s}},\nu)B(\nu)I(\hat{\mathbf{s}},\nu)e^{-i2\pi\nu\mathbf{U}.\hat{\mathbf{s}}}d\Omega,$ (4.4)
@@ 5 Emulating Hi 21cm Power Spectrum
Figure 3: The figure shows a comparison between the simulated true power spectrum (solid lines) and the emulated power spectrum by the ANN (dots). The left panel shows emulator predictions trained on theoretical power spectra, while the right panel presents predictions based on observed power spectra from SKA-Low.
@@ 7 Error Covariances of Power Spectrum
where $\rm A_{\mathrm{eff}}$, $\rm A_{\mathrm{coll}}$, and $\rm A_{\mathrm{core}}$ denote the effective area, collecting area, and core area, respectively, of the specified radio interferometric array. The symbol $\rm t_{\mathrm{obs}}$ signifies the total observation time in hours. Additionally, $\rm D$ and $\rm\Delta D$ denote the comoving distance to the redshifts where the center is located and the comoving distance corresponding to a bandwidth $\rm B$ at that comoving distance. $\rm T_{sky}$ denotes the sky temperature calculated using $\mathrm{T_{\mathrm{sky}}\sim 180\Big(\frac{\nu}{180MHz}\Big)^{-2.6}}$K [5], symbol $\Omega_{\mathrm{FoV}}$ denotes field of view symbol, $n_{\mathrm{p}}$ represents the number of polarization and $N_{\mathrm{k_{i}}}$ denotes the number of measurement at each k-mode.
@@ 8.1 Perfect Observing Condition
We also noticed that the power spectrum from set 1, whose astrophysical parameters are around the centre of the boundaries, generated more accurate conclusions. By contrast, the power spectra generated in close proximity to the boundary values of the astrophysical parameters yielded an inferred parameter mean that slightly differed from the actual value. For example, in Set 3, the true value of $\rm\zeta$ is 57.42, while the inferred mean using the theoretical power spectrum is 50.8. Using the SKA power spectrum, the inferred mean is 59.8. Although these values differ from the true mean, they still lie within $1\sigma$ of the inferred uncertainty range (see Table 2). This mismatch could be due to the ANN emulator being trained on less training datasets. Potential future development of a more precise emulation could involve training it with a more extensive dataset that cover the parameter space in the effective manner.
@@ 8.2.1 Gain Calibration Errors
To assess the influence of gain calibration errors on the observed power spectrum (PS) and their subsequent impact on associated astrophysical parameters for SKA-Low, we introduced various gain calibration errors as detailed in [1]. The resulting residual PS was visually compared with the actual PS, depicted in the left panel of Fig. 5. The parameter obtained from the PS analysis with SKA-Low, incorporating actual signal and PS for gain error residuals, is depicted in the right panel of Fig. 5. Upon examining the plotted residual PS spectrum, we observed that for gain calibration error of 0.001 %, the residual PS overlaps with the true PS. For residual PS with a gain calibration error of 0.01%, we observe overlapping at lower k modes but with minor deviation at higher k modes. However, when deriving astrophysical parameters from the residual power spectrum, we found that for a gain calibration error of 0.001%, the derived parameters match those derived from the true PS. In contrast, for PS with the gain calibration error of 0.01%, the derived parameters exhibit significant differences from the actual values. This indicates that astrophysical parameters are highly sensitive to gain calibration errors, and even minor deviations in the PS due to gain calibration errors can severely impact the estimation of parameters, introducing bias. The sensitivity of derived parameters to gain calibration errors clearly suggests that the gain calibration error threshold should be much lower than 0.01%, ideally closer to 0.001%. If the gain calibration error exceeds 0.001%, it is advisable to model them to prevent bias in the derived parameter values.
@@ 8.2.2 Sky Model Position Error
Similar to gain calibration errors, we evaluated the impact of sky model position errors on the observed power spectrum (PS) and their effect on the inferred astrophysical parameters for SKA-Low. Different sky model position errors, as described above, were introduced, and the resulting residual PS was compared with the true PS, as shown in the left panel of Fig. 6. The parameters obtained from the PS, which incorporates actual signals and position error residuals, are depicted in the right panel of Fig. 6. Upon examining the plotted residual PS spectrum, we found that for a sky model position error of 0.048”, the residual PS overlaps with the true PS. At lower k modes, the residual PS with this position error overlaps well with the true PS, but at higher k modes, there is a minor deviation. This occurs because real-space scales are more affected by arcsecond-level errors at higher k values. However, since we used the entire k range to derive astrophysical parameters from the residual power spectrum, we observed that for a position error of 0.048”, the derived parameters closely match those obtained from the true PS. Similarly, for a sky model position error of 0.48”, the derived parameters exhibit substantial deviations from the true values. This indicates that astrophysical parameters are highly sensitive to sky model position errors, and even small deviations in the PS due to these errors can severely impact parameter estimation, introducing bias. The sensitivity of derived parameters to sky model position errors, similar to gain calibration errors, suggests that the acceptable sky model position error threshold should be much lower than 0.48”, ideally around 0.048”. If the sky model position error exceeds 0.048”, modeling the errors becomes crucial to prevent bias in the derived parameter values.
@@ 9 Summary and Discussion
This study presents the development of an ANN and Bayesian-based framework for inferring astrophysical parameters from observed Hi power spectra (PS). We used the 21cm E2E pipeline to generate the mock observed power spectrum (PS) for SKA-Low core array configurations. However, generating a large number of modeled observable signals across a multi-dimensional parameter space for Bayesian inference is computationally expensive, particularly when using simulated observations. To address this, we adopted a formalism utilized by several groups, employing emulators for modelling the EoR signal instead of actual simulations.
@@
  • The position error threshold for SKA-Low is found to be <0.48<0.48", as displacements >0.048>0.048" cause significant deviations between the inferred astrophysical parameters and their actual values. This suggests that sky model position errors can bias the inference of the astrophysical parameters.
@@
Thus, we conclude that for the upcoming SKA Low interferometer, the gain calibration error should be nearly 0.001%, and the sky model position error should be less than 0.48” to avoid biasing the inference. Beyond these thresholds, these errors start to affect the results. To reduce their impact, calibration must be performed with the same level of accuracy, or we need to develop modeling or mitigation techniques to eliminate the residual foregrounds arising from calibration or sky model inaccuracies, ensuring minimal influence on the inference. It is important to note that the simulations presented in this study rely on several simplifying assumptions regarding foregrounds and instrumental effects. This underscores the need for more realistic and detailed simulations, including factors such as primary beam chromaticity and other instrumental systematics, to more accurately assess the robustness of parameter inference and to inform optimal array design. These complexities will be addressed in future work for a more comprehensive understanding.
Acknowledgments